# The Battle of the Water Futures

### v1

@@
Dennis Zanutto    Christos Michalopoulos    Lydia Tsiami    Andre Artelt    Jasmin Brandt    Demetrios Eliades    Stelios Vrachimis    Stefano Alvisi    Valentina Marsili    Filippo Mazzoni    Panagiotis Samartzis    Barbara Hammer    Phoebe Koundouri    Marios Polycarpou    Dragan Savić
@@ 1.1 Disclaimer
The networks and problem presented in this competition are synthetic and do not represent any real-world situation. While we utilize real data from trusted sources where possible (e.g., CBS for municipal data, KNMI for wheater observations), other aspects rely on assumptions based on expert knowledge.
@@
All versions are archived on Zenodo, with the latest version establishing the official standard. An open-source evaluator will also be released in the WaterBenchmarkHub (Artelt, 2025), allowing competitors to test their solutions before the final submission.
@@ 1.2 Background and motivation
Water distribution systems (WDS) grow organically with cities and are therefore inherently subject to deep uncertainties. While secondary and tertiary network components can still be adequately selected using established engineering principles, planning the primary and supply network at a regional level requires a different approch that integrates elements of policy-making, standard engineering practice and considerable investments. This part of the system is extremely critical as failures here can propogate and lead to loss of service for thousands of people and businesses. Beyond system reliability and deep uncertainties, the complex governance requirements, such as the coordination between national, regional, and local administrations, present a major obstacle to integrated policy-making and long-term planning of these systems.
Over the past decade, the field has steadily refined and built upon several approaches that address WDS planning under uncertainties. To begin, the introduction of staged design, where interventions are modeled in phases rather then in as single batch representing the target network, provided a tool to represent long-term plans with more realism (for an example on WDS see Creaco et al., 2014). Next, this modeling framework was extended to include uncertainty, either through a robust approach that accounts for uncertainty (e.g., Creaco et al., 2015), or through a flexible approach where one plan with multiple pathways (in its simplest form: a decision tree) is devised, but only one branch is implemented based on the uncertainties realisation (e.g., Basupi and Kapelan, 2015).
An inherent limitation of robust and flexible plans is that they require accounting for all possible future system interventions and scenarios at the initial planning stage. Therefore, when uncertainties move away from the accounted values or new opportunities emerge (i.e., under deep uncertainties), these approaches fall short: the water utilities must adapt to the changes in their environment.
An adaptive plan would typically feature at least one of two components. The first is a periodic re-evaluation and adjustment of the plan based on uncertainty evolution to incorporate new opportunities and information (see, for an example, Beh et al., (2017) on water supply systems, and Skerker et al., (2023) on “the value of learning” in flexible planning for water reservoirs). The second is the ability to shift between pathways and avoiding locking-up to one (see dynamic adaptive policy pathways for the Rhine Delta, Haasnoot et al., 2013).
@@
Therefore, in the context of the Battle of the Water Futures, the participants’ goal is to apply novel methodologies to develop masterplans for a nation’s water utilities under deep uncertainties; while the orgainsers objective is to provide a benchmark that advances the state of practice in decision-making under deep uncertainty for water distribution systems and establishes a foundation for benchmarking through a highly customizable, open-source framework.
@@ 1.3 Ambition
Participants must develop masterplans for a nation’s water utilities operating across different provinces. However, the BWF presents a level of detail and computational complexity such that the featured transport network optimization problem cannot be solved through brute force optimization. Moreover, unlike traditional optimization problems with specified scenarios, here, only a description of the environment dynamics and a handful of expert-driven scenarios for key drivers (e.g., population) are provided. Competitors must embrace uncertainty and decide for themselves what to model and how, reflecting the reality that different teams would make different modeling choices when facing deep uncertainty.
The decision to place participants in a situation of deep uncertainty is deliberate. While some parameters are uncertain and the evolution of these uncertainty bounds in the future will not be disclosed (e.g., the annual cost of operating a production facility, which varies between minimum and maximum values), past observations will be made available at the end of each competition stage, enabling participants to implement adaptive strategies and learn from accumulated evidence. Similarly, the competition ranking process has not been announced, only the system requirements are explained in this document (i.e., the metrics under which the system is evaluated), compelling participants to identify their own compromise solutions within this many-objective context.
@@ 2 Problem Description
The Battle of the Water Futures takes place in a not so ficticious world, with a blend of real data and assumptions. The BWF is concerned with the optimal planning and operation of multiple regional water distribution systems in a national landscape. It is based on The Netherlands where KWR is located and instead of the current water ulitities division, one independent water utility is assumed for every province. Each water utility will need to prepare a masterplan for at least the next 25 years and run the operations of their systems. Competitors will also make some decision regarding national and regional policies, like budget allocation and price increase. Various information about the current status of the system (network model, prices, demands, population, etc) are made available, consult DATA AVAILABLE page. Together wit hthe current status, future projections of the relevant drivers (e.g., population, electtricity shares) can be used to formulate the plan. The masterplan consist of a series of interventions, policy decisions and operations that can be applied over the years of the planning horizon, this effectively requires staged design and applied phased interventions.
Once the masterplan is completed, the utilities will start following it and operate and mantain the system accordingly to the specifications. The world will evolve, uncertainties will be realized (shown) and, hopefully, they will be able to fullfill their primary objective: satisfying the consumers water demand with adequate level of service.
During the time when the masterplan is being executed, several key performance indicator are extracted, from the main finance health and reliability (level of service), to secondary objectives like GHG emissions and fairness. The water utilities will also have access to data about their networks, such as pressures, demands, power consumptions etc
After 25 years, the utilities will have a chance to update their masterplan, adapting it based on the new acquired informaiton (now history), and the cycle will repeat (plan, execute, observe).
Both interventions and input parameters can be assumed to be deep uncertaintites, meaning that only approximate definition of their future evolution are provided and that no probability (or wrong ones) are associated with the realisation of each scenario. Everything should be regarded only as an indication, to the best of the current knowledge, made by some experts. Take these as an example: - the time to construct a desalination plant: it can be defined between 5 and 10 years. Once construction starts its actual implementation time can be anything between those bounds - population evolution: projections of population in the country wille be provided with a mean value and upper and lower bounds with various uncertainty level. However, how this parameter will evolve no one can know Behind the scenes we blend experts opinions with a dynamic model with tunable parameters to to generate the synthetic time serieses that will be the input of this system.
Planning horizon: free, but at least 25 years Interventions resolution (masterplan intervnetions timestep): year/quarter Stage horizon/masterplan update period: 25 years Hydraulic simulation timestep: 1 hour Every other parameter will have its own resolution as defined in the input files. Simulation approach: every day for the whole stage > if too computationally demanding (>20 minute per solution), we reduce the number of days mantaining the same timestep, i.e., 1 week representative of each month or 1 week representative of each quarter/season.
Definitions Deep uncertainties Staged Design Robust planning Flexible planning Adpative planning
2.1 System Decription
@@
This chapter describes the physical components that constitute the water transport network in the BWF competition. We present the key entities—municipalities, water sources, pumping stations, and pipes—detailing their characteristics, operational constraints, and how they evolve over time. Each component’s properties, costs, and decision variables are outlined to provide participants with a complete understanding of the system they will manage. The physical feasibility of any network configuration is verified through hydraulic simulation in EPANET. External factors affecting the system, such as climate conditions and the electricity grid, are addressed separately in Section 2.2.
@@ 2.1.1 Water Utilities
WUs must chose the policies to apply to their own system (e.g, setting the budget for non-revenue water reduction) and decide the interventions to take on their system (e.g., replacing a pipe).
Water utilities can also join their efforts to pursue a common interventions, for example, the cost of placing a pipe connecting two water utilities will be shared between the two utilities.
The structure of the system is static. The number of WUs and their specific geographic responsabilities (assigned provinces) will not change through the competition.
@@
Property Type Scope
@@ 2.1.2 Municipalities
Along with these attributes, the only observable quantity for each municipality is the total water demand divided in its two components: consumption (delivered demand) and the quota of undelivered demand.
More precisely, the total water demand comprises two volumetric quantities, though this breakdown is not observable to participants:
  • billable water demand (the sum of household and business water demands) described in Section 2.1.2.1, and
  • non-revenue water (accounting for leaks, flushing, measurement errors and other losses), described in Section 2.1.2.2.
Instead, only two outputs are observable: actual consumption and undelivered demand. These two variables are extracted from an EPANET simulation of the network run in pressure-driven analysis (PDA) mode with a minimum pressure threshold of 30 m. This threshold accounts for the requirement to deliver 20 m of pressure at the consumer tap, plus an allowance of 5 m pressure drop in the secondary network and 5 m in the tertiary portion.
@@ 2.1.2.1 Water Demand Model
The methodology developed to generate water consumption time series builds on historical data from the Dutch Drinking Water Statistics(Vewin 2025) , which provide nationwide trends in total drinking water production, sectoral water use, and non-revenue water over the period 2000–2024. Specifically, water-consumption time series generation is structured into three phases.
Phase I. The first phase estimates the annual water volume supplied to each municipality using information on households, and businesses, complemented by projected data where required. These annual volumes are calibrated to match national totals reported in official statistics and then randomized around the calibrated value to introduce variability among municipalities.
@@
Phase III. The third phase produces the final hourly time series by applying a Fourier series-based approach which combines seasonal modulation, climate-related adjustments, and random perturbations to capture temporal variability. The two residential profiles associated with each municipality are aggregated through weighted combinations, and both residential and non-residential profiles are scaled to match the previously estimated yearly volumes. The outcome is a set of 8,760-point hourly water consumption series for each municipality.
@@
Daily per household demand Dynamic Exogenous Municipality L/house/day
Daily per business demand Dynamic Exogenous Municipality L/business/day
Max yearly temperature Dynamic Exogenous National [°C]
@@ 2.1.2.2 Non-Revenue Water Model
Our modelling of non-revenue water (NRW) relies on the average age of the pipe infrastructure in a municipality. Based on this average age, the municipality gets assigned to one of five possible classes (A - E) as follows:
Age in years Class Lower and upper bound of NRW demand in mˆ3/day
0 - 25 A (0, 12)
25 - 43 B (12, 20)
43 - 54 C (20, 35)
54 - 60 D (35, 55)
> 60 E (55, inf)
Each class is associated with a distribution of non-revenue water (NRW)demands (e.g., leakages), from which we sample to generate the total demand, consisting of the normal demand + NRW demand. Notably, the NRW demand is different for each class – i.e., older systems suffer from more leaks and therefore have a higher NRW demand. The distribution of NRW demands per class and km of pipes is illustrated in the following figure.
Figure 1: leak-demands
The total number of km of pipes in a given municipality is linked to its population size. Here, we use a linear relationship between the population size and the km of pipes, also illustrated in the following Figure:
km_pipes=57.7population_in_10kkm\_pipes=57.7*population\_in\_10k
Figure 2: pop-pipes
@@
Property Type Scope
NRW demand Static [Uncertain] National
Inner distribution network - length to population ratio Static National
Inner distribution network - length Dynamic endogenous Municipality
Inner distribution network - average age Dynamic endogenous Municipality
NRW intervention - unit cost Dynamic endogenous NRWClass, Municipality Size Class, National
@@
Intervention budget Option

### v4

@@
Dennis Zanutto    Christos Michalopoulos    Lydia Tsiami    André Artelt    Jasmin Brandt    Demetrios Eliades    Stelios Vrachimis    Stefano Alvisi    Valentina Marsili    Filippo Mazzoni    Panagiotis Samartzis    Barbara Hammer    Phoebe Koundouri    Marios Polycarpou    Dragan Savić
@@
Executive Summary
This competition challenges participants to confront one of water distribution systems’s most pressing problems: adaptive planning under deep uncertainty. Participants will develop a 75-year national masterplan for a hypothetical Dutch water grid that must balance four competing objectives: financial sustainability, service reliability, environmental impact, and equitable access. Unlike traditional planning exercises, this competition requires strategic decisions across three 25-year stages while navigating uncertain climate, energy, and economic scenarios. Success demands developing an adaptive masterplan that blends policy decisions and infrastructure interventions, gets updated based on evidence accumulated across the implementation stages, and preserves flexibility to avoid locking into suboptimal pathways. This document defines the competition framework; a subsequent report will present participant solutions and outcomes.
@@ 1.1 Disclaimer
The networks and the problem presented in this competition are synthetic and, despite being based on the Netherlands, do not represent any real-world situation. While we utilize real data from trusted sources where possible (e.g., CBS [^1] CBS (Centraal Bureau voor de Statistiek) is the Dutch national statistics agency: https://www.cbs.nl for municipal data, KNMI [^2] KNMI (Koninklijk Nederlands Meteorologisch Instituut) is the Dutch meteorological institute: https://www.knmi.nl for weather observations, Vewin [^3] Vewin is the association of Dutch drinking water companies: https://www.vewin.nl for statistics about water utilities), other aspects rely on assumptions based on expert knowledge.
@@
All versions are archived on Zenodo, with the latest version establishing the official standard (Zanutto et al. 2025). An open-source evaluator will also be released in the WaterBenchmarkHub (Artelt et al. 2025), allowing competitors to test their solutions before the final submission.
@@ 1.2 Background and motivation
Water distribution systems (WDS) grow organically with cities and are therefore inherently subject to deep uncertainties. Strategic WDS planning focuses on the long-term design of the primary supply system – the national grid. While secondary and tertiary network components can typically be selected using established engineering principles, planning the primary and supply network at a regional level requires a different approach, one that integrates policy considerations, standard engineering practice, and substantial investment. This part of the system is extremely critical. Failures at this level can propagate and lead to loss of service for thousands of people and businesses. Beyond system reliability and deep uncertainties, the complex governance requirements, such as the coordination between national, regional, and local administrations, present a major obstacle to integrated policy-making and long-term planning of these systems.
Over the past decade, the field has steadily refined and built upon several approaches that address WDS planning under uncertainties. To begin, the introduction of staged design, where interventions are modelled in phases rather than in a single batch representing the target network, provided a tool to represent masterplans with more realism (e.g., Creaco, Franchini, and Walski 2014). Next, this modelling framework was extended to include uncertainty, either through a robust approach that accounts for it (e.g., Creaco, Franchini, and Walski 2015), or through a flexible approach that develops a single plan with multiple pathways. In its simplest form, this takes the shape of a decision tree, where only one branch is implemented based on how uncertainties unfold (e.g., Basupi and Kapelan 2015).
An inherent limitation of robust and flexible plans is that they require accounting for all possible future system interventions and scenarios at the initial planning stage. Therefore, when uncertainties move away from the expected values or new opportunities emerge (i.e., under deep uncertainties), these approaches fall short, and the water utilities must adapt to the changes in their environment.
An adaptive plan would typically feature at least one of two components. The first component involves periodically re-evaluating and adjusting the plan as uncertainties evolve, allowing new opportunities and information to be incorporated. For example, see Beh, Maier, and Dandy (2015) on water supply systems, and Skerker et al. (2023) on “the value of learning” in flexible planning for water reservoirs. The second is the ability to shift between pathways and avoid locking up to one (see dynamic adaptive policy pathways for the Rhine Delta, Haasnoot et al. 2013).
@@
Therefore, in the context of this Battle of the Water Futures (BWF), the participants’ goal is to apply innovative methodologies to develop a masterplan for a national water grid under deep uncertainties. The organisers, in turn, seek to establish a benchmark that advances decision-making practices for water distribution systems under deep uncertainty, while laying the ground for future benchmarking through a highly customizable, open-source framework.
@@ 1.3 Ambition
Participants must develop masterplans for a national water grid with water utilities operating across different provinces. However, the BWF presents a level of detail and computational complexity such that the featured national grid optimisation problem cannot be solved through brute force optimisation. Moreover, unlike traditional optimisation problems with specified scenarios, here, only a description of the environment dynamics and a handful of expert-driven scenarios for key drivers (e.g., population) are provided. Competitors must embrace uncertainty and decide for themselves what to model and how to do so, reflecting the reality that different teams would make different modelling choices when facing deep uncertainty.
The decision to place participants in a context of deep uncertainty is intentional. The BWF will unfold over three competition stages, each spanning 25 years, to realistically emulate the long-term challenges faced by planners. This means that some parameters are uncertain and their uncertainty bounds evolution will be provided only through approximate expert-based scenarios (e.g., population projections vary between minimum and maximum values, though actual outcomes may fall outside these bounds). At the end of each competition stage, past observations will be made available, enabling participants to implement adaptive strategies and learn from accumulated evidence. Similarly, the competition ranking process has been deliberately left undisclosed; only the system requirements, namely, the metrics under which the system is evaluated, are outlined in this document. This approach is more akin to realistic situations and compels participants to identify their own compromise solutions within this many-objective context.
@@ 2 Problem Description
The Battle of the Water Futures (BWF) problem focuses on a hypothetical national drinking water grid of the Netherlands. The system structure (demand nodes, sources, and links), available interventions (e.g., pipes, treatment plants), and exogenous drivers (e.g., population, climate) are based on publicly available data about the Dutch context, supplemented by assumptions or insights from other regions as needed.
The main objective of the BWF is to keep the system operational until the end of the century, thereby preventing bankruptcy while fulfilling the required service requirements in each water utility. Secondary objectives include reducing emissions toward climate neutrality and ensuring fairness across regions and generations. To achieve these goals, the country’s water utilities come together to develop a national masterplan, which they agree to update every 25 years.
The masterplan specifies the interventions to be applied each year, both at the national level (e.g., connecting independent water utilities) and at the individual utility level (e.g., opening new sources). The masterplan does not need to include the replacement of ageing components (pipes and pumps), as these are automatically replaced when they reach their end of life. However, to manage this uncertainty, utilities can either schedule replacements in advance or allow components to age naturally and account for replacement costs in the budget. The masterplan also defines the national and regional policies and whether they are revised during the planning period (note: while interventions must be specified annually, policies can be set initially and remain as they are until amended in the plan).
To prepare the masterplan, information on the current system status is made available, including network models, prices, demands, population, and other relevant parameters. Future projections of key drivers (e.g., population growth, climate trends) are also provided to complement system knowledge. However, these scenarios are based on current expert knowledge and should not be treated as precise forecasts as these variables are generally deep uncertainties of the system.
Examples of deep uncertainties include:
  • Desalination plant construction time: Planners are confident it can be built in 5 years, but acknowledge that delays could extend completion to 10 years.
  • Population evolution: Population is expected to follow the mean values or remain within the upper and lower bounds described by the national statistics agency. However, major external events, such as conflicts or mass migration driven by climate catastrophes, could push evolution beyond these ranges.
Behind the scenes, expert opinions are combined with dynamic models to generate the synthetic time series that drive system evolution. At each timestep, EPANET hydraulic simulations verify the network’s physical capability to transfer water and meet demands under the proposed interventions and operating conditions.
Once the masterplan is defined, utilities begin following it, operating and maintaining their systems according to its specifications. Around them, the world evolves and uncertainties unfold. After 25 years, utilities reconvene to update their plan. All system information is updated, and the performance over the past 25 years along with observed variables can be used to revise or completely redraw the masterplan. This cycle—plan, execute, observe, adapt—repeats for three rounds, testing the utilities’ ability to manage both short-term operations and long-term strategic planning under deep uncertainty.
2.1 System Description
@@
This chapter describes the physical components that constitute the water transport network in the BWF competition. The key entities are municipalities, water sources, pumping stations, and pipes and each of the following section will detail their characteristics, operational constraints, and how they evolve over time. The physical feasibility of any network configuration is verified through hydraulic simulation in EPANET. External factors affecting the system, such as climate conditions and the electricity grid, are addressed separately in Section 2.2.
@@ 2.1.1 Water Utilities
WUs must choose the policies to apply to their own system (e.g, setting the budget for non-revenue water reduction) and decide the interventions to take on their system (e.g., replacing a pipe).
Multiple WUs can also join their efforts to pursue common interventions, for example, the cost of placing a pipe connecting two water utilities will be shared between the two utilities. If WU are connected with each other and there is an exchange of water, the receiving utility pays a volumetric fee to the supplying utility. More detailed information about the water tariff scheme are available in Section 2.2.3.3.
The structure of the system is static. The number of WUs and their specific geographic responsibilities (assigned provinces) will not change through the competition.
@@
Property Type Scope Unit
@@ 2.1.2 Municipalities
Number of businesses Dynamic Exogenous Municipality units
Average Disposable Income Dynamic Exogenous Municipality $k\text{\texteuro}$
@@
The total municipality water demand comprises two volumetric quantities:
  • Billable water demand ($D^{\text{BIL}}$): The sum of household and business water demands described in Section $2.1.$
  • Non-revenue water ($D^{\text{NRW}}$): accounting for leaks, flushing, measurement errors and other losses described in Section 2.1.2.2.
While these quantities represent the physical components of demand, they are not directly observable by participants. Instead, participants observe the total water demand divided into two components: consumption ($Q$; delivered outflow) and undelivered demand ($U$).
For each municipality $m$ at time $t$, these quantities maintain the following relationships:
$D_{m}(t)$ $=D^{\text{BIL}}_{m}(t)+D^{\text{NRW}}_{m}(t)$ (1)
$=Q_{m}(t)+U_{m}(t)$
The decomposition between delivered and undelivered demand is extracted from an EPANET simulation of the network run in pressure-driven analysis (PDA) mode with a minimum pressure threshold of 30 m. Whenever there is undelivered demand, we assume that this reduces the billable component first, i.e., $Q^{\text{BIL}}_{m}(t)=D^{\text{BIL}}_{m}(t)-U_{m}(t)$.
@@ 2.1.2.1 Water Demand Model
The methodology developed to generate water consumption time series builds on historical data from the Dutch association of water companies (Vewin 2025), which provide nationwide trends in total drinking water production, sectoral water use, and non-revenue water over the period 2000–2024. Specifically, water-consumption time series generation is structured into three phases.
Phase I. The first phase estimates the annual water volume supplied to each municipality using information on households and businesses (CBS 2025), complemented by projected data where required. These annual volumes are calibrated to match national totals reported in official statistics (Vewin 2025) and then randomized around the calibrated value to introduce variability among municipalities.
@@
Phase III. The third phase produces the final hourly time series by applying a Fourier series-based approach which combines seasonal modulation, climate-related adjustments (accounting for the maximum yearly temperature), and random perturbations to capture temporal variability. The two residential profiles associated with each municipality are aggregated through weighted combinations, and both residential and non-residential profiles are scaled to match the previously estimated yearly volumes.
Therefore, the total billable demand of municipality $m$ at time $t$ (within year $y$) is defined as:
[…]
where $D^{\text{R1}}_{m}(t,T_{y})$ and $D^{\text{R2}}_{m}(t,T_{y})$ represent the two residential demands, $w_{m}\in[0,1]$ is the unitary weight to combine them, $D^{\text{C}}_{m}(t,T_{y})$ is the (commercial) non-residential demand, and $T_{y}$ is the maximum temperature recorded in year $y$.
@@
Daily per household demand Dynamic Exogenous Municipality m³/house/hour
Daily per business demand Dynamic Exogenous Municipality m³/business/hour
Max yearly temperature Dynamic Exogenous National °C
@@ 2.1.2.2 Non-Revenue Water Model
Non-revenue water (NRW) is an uncertain quantity modeled through the average age of pipe infrastructure in each municipality’s inner distribution network (IDN). Based on this average age, municipalities are assigned to one of five NRW classes as reported in Table LABEL:tbl:nrw-classes. Each class is associated with a distinct probability distribution of NRW demands, from which daily samples are drawn to generate the volumetric NRW demand factor ($m^{3}/km/day$). Notably, older infrastructure suffers from more leaks and therefore exhibits higher NRW demand factors.
The distribution of NRW demands varies by class and is illustrated in Figure 1. The actual values of the distributions’ parameters can be inspected within the data files, which are mapped in Appendix A.
The total length of pipes in a municipality is linked to its population size through the following linear relationship:
[…]
where $m$ is the municipality index, $y$ is the year, and $L^{\text{IDN}}_{m}(y)$ is the total length of pipes (km) in municipality $m$ at year $y$.
The actual municipality NRW demand is also capped at twice the billable daily demand to prevent unrealistic leakage levels. Therefore, total NRW demand for municipality $m$ at day $d$ is calculated as:
[…]
where $f^{\text{NRW}}_{\text{class}(m)}$ is the sampled NRW demand factor ($m^{3}/km/day$) for the municipality’s class, and $\bar{D}^{\text{BIL}}_{m}(d)$ is the average daily water demand of municipality $m$ at day $d$. The daily leak $D^{\text{NRW}}_{m}(d)$ is equally spread across the day (i.e., $D^{\text{NRW}}_{m}(t)=D^{\text{NRW}}_{m}(d)/24$).
Figure 1: Non-revenue water demand factor per class
Table 4: Non-revenue water classification by infrastructure age.
Inner distribution network - average age [years] NRW Class Probability distribution
0 - 25 A Inverted Exponential
25 - 43 B Uniform
43 - 54 C Uniform
54 - 60 D Uniform
> 60 E Exponential
@@
Property Type Scope Unit
Inner distribution network - length to population ratio Static National $km\cdot(10^{4}\text{ inhabitants})^{-1}$
Inner distribution network - length Dynamic endogenous Municipality kmkm
Inner distribution network - average age Dynamic endogenous Municipality yearsyears
NRW intervention - unit cost Dynamic endogenous NRWClass, Municipality Size Class, National $\text{\texteuro}/year/km$
@@
Intervention budget Option $\text{\texteuro}/year$